Help 6 The following is a report from a BLS survey taker in the houses I visited. Ten of them were children under 16, 25 people had full-time jobs, and 5 had part-time jobs. There were 10 retirees, 5 full-time homemakers, 5 full-time students over age 16, and 2 people who were disabled and cannot work. The remaining people did not have jobs, but all said they would like one. One of these people had not looked actively for work for 3 months, however .5 Find the labor force, the unemployment rate, and the participation rate implied by the survey takers report. Labor force: 32 people Instructions: Enter your responses rounded to two decimal places Unemployment rate: 62 % Participation rate: [mm-Too
Answer #1

Only individuals age 16 and older are counted in the potential labor force, and they are only counted in the labor force if they are actively working or have looked for work in the past four weeks.

Remaining people = 65-10-10-25-5-5-5-2= 3

Labor force = 25 full time + 5 part time+ 3 remaining -1 not looking for work = 32

Labor force = 32

No. of Unemployed people = 3-1= 2

Unemployment rate = (No. of unemployed / labor force) x 100 = 2/32*100= 6.25%

Working-age population (ages 15 to 64 in most OECD countries)= (65-10-10) = 45 [ here 10 students under 16 and 10 retirees]

Labor Force Participation Rate = Labor Force/ Working Age Population = 32/45= 0.71